OpenStudy (anonymous):

Find the maximum and minimum of the values of f on the given interval and sate where these values occur: f(x)= x^2/3 (20-x); [-1, 20]

OpenStudy (misssunshinexxoxo):

This might help http://cnx.org/content/m17417/latest/

OpenStudy (superdavesuper):

max n min occur when f'(x)=0 so u need to differentiate the original eqn to find f'(x).

OpenStudy (anonymous):

Yeah I know how to do it I'm just stuck on one part

OpenStudy (superdavesuper):

which part? :)

OpenStudy (anonymous):

Umm finding the critical points

OpenStudy (superdavesuper):

What is f'(x)? u can find critical pts by setting f'(x)=0

OpenStudy (misssunshinexxoxo):

Dave maybe solve the equation and offer evidence to support

OpenStudy (superdavesuper):

Okiee...not sure if I am supposed to but f(x)=20x^2/3 -x^5/3 So f'(x)=20*2/3*x^(-1/3) - 5/3*x^2/3 Then set f'(x)=0 20*2/3*x^(-1/3) - 5/3*x^2/3=0

OpenStudy (superdavesuper):

Is this where u are stuck?

OpenStudy (anonymous):

Yep :(

OpenStudy (superdavesuper):

divide everything by 5/3 first...

OpenStudy (anonymous):

Okay I get 8x^-1/3-x^-1/3

OpenStudy (superdavesuper):

Yup yup then re-arrange by taking x^(-1/3) out so you get x^(-1/3) * (.......) what is inside the bracket?

OpenStudy (anonymous):

Sorry I ment * x^2/3

OpenStudy (superdavesuper):

yup yup saw that - no prob at all :) u are doing just fine....almost there!

OpenStudy (anonymous):

Umm I'm not sure how to factor it out...